• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name septum_auto

    Log entry time 23:23:38 on May 2, 2012

    Entry number 381563

    keyword=Wrong septum current, please check!

    Left septum set current 362A is not correct, should be 725.4A! please check!

    Right septum set current 475A is not correct, should be 919.752A! please check!

    septum alarm using halog_post [halog_post 1.1]

    A copy of this log entry has been emailed to: jixie@jlab.org